OpenStudy (anonymous):

A set of 50 data values has a mean of 22 and a variance of 16. I. Find the standard score (z) for a data value = 14. II. Find the probability of a data value < 14.

OpenStudy (anonymous):

1).a set of 50 data values has a mean of 40 and a variance of 25. a). Find the standard score (z) for a data value = 47. z(47) = (47-40)/25 = 0.28 b). Find the probability of a data value > 47. P(x> 47) = P(z > 0.28) = 0.3897 (2). Find the area under the standard normal curve: I. to the right of z = 1.83 P(z > 1.83) = 0.0336 II. to the left of z = 1.83 P(z < 1.83) = 1-0.0336 = 0.9664

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

bradley Your Welcome :D